6.What number when divided by one half is twelve less than three times the number?
F 24
G -12
H 12
J 18

Answer by checkley77(12569)

X/.5=3X-12
X/.5-3X=-12
2X-3X=-12
-X=-12
X=12 ANS. H).
PROOF:
12/.5=3*12-12
24=36-12
24=24